# VLMs Play StarCraft II: A Benchmark and Multimodal Decision Method

### v2

@@ Abstract
We introduce VLM-Attention, a multimodal StarCraft II environment that aligns artificial agent perception with the human gameplay experience. Traditional frameworks such as SMAC rely on abstract state representations that diverge significantly from human perception, limiting the ecological validity of agent behavior. Our environment addresses this limitation by incorporating RGB visual inputs and natural language observations that more closely simulate human cognitive processes during gameplay. The VLM-Attention framework consists of three integrated components: (1) a vision-language model enhanced with specialized self-attention mechanisms for strategic unit targeting and battlefield assessment, (2) a retrieval-augmented generation system that leverages domain-specific StarCraft II knowledge to inform tactical decisions, and (3) a dynamic role-based task distribution system that enables coordinated multi-agent behavior. Our experimental evaluation across 21 custom scenarios demonstrates that VLM-based agents powered by foundation models (specifically Qwen-VL and GPT-4o) can execute complex tactical maneuvers without explicit training, achieving comparable performance to traditional MARL methods that require substantial training iterations. This work establishes a foundation for developing human-aligned StarCraft II agents and advances the broader research agenda of multimodal game AI. Our implementation is available at https://github.com/camel-ai/VLM-Play-StarCraft2.
@@ 1 Introduction
We present VLM-Attention, a novel StarCraft II environment that fundamentally reimagines how AI agents perceive and interact with the game world. Our framework makes three key innovations: First, we redesign the observation space to align with human cognitive processes, incorporating both RGB visual input and natural language descriptions. This multimodal approach enables agents to process battlefield information in a manner analogous to human players, facilitating more intuitive strategic decision-making. Second, we develop a comprehensive action space that supports complex military maneuvers, including unit targeting, formation control, and ability usage. This design allows for fine-grained tactical control while maintaining accessibility for high-level strategic planning. Third, we introduce an integrated agent architecture that combines the reasoning capabilities of VLMs with specialized mechanisms for multi-agent coordination. We incorporate self-attention mechanisms for strategic unit targeting, retrieval-augmented generation (RAG) for leveraging domain-specific knowledge, and dynamic role assignment for effective task distribution.
To validate our approach, we create a diverse set of 12 micromanagement scenarios that test various aspects of tactical decision-making and unit control. Our experimental results demonstrate that VLM-based agents, utilizing models such as Qwen-VL and GPT-4o, can effectively execute complex tactical maneuvers without requiring extensive training, a significant departure from traditional MARL approaches. The primary contributions of this work include:
  • A novel multimodal StarCraft II environment that aligns agent perception with human cognitive processes, featuring RGB visual input, natural language observations, and a comprehensive action space for complex unit control.
  • An integrated VLM-based agent architecture that combines self-attention mechanisms, retrieval-augmented generation, and dynamic role assignment for effective tactical decision-making.
  • Empirical evidence demonstrating the effectiveness of VLM-based agents in executing complex StarCraft II micromanagement tasks without explicit training, supported by extensive evaluation across 12 scenarios.
Our work establishes a foundation for developing more human-aligned StarCraft II agents and contributes to the broader goal of creating AI systems that can understand and interact with complex game environments in ways that more closely mirror human cognition. In the supplementary material, we provide the code implementation and video demos of our agent decision-making.
@@ 2 Related Work
3 Environment Design
Figure 1: Observation Space of environment.
Table 1: Comparison of StarCraft II environments. While SMAC simplifies the interface for MARL training and PySC2 provides comprehensive API access, VLM-Attention adapts PySC2 for multimodal decision-making through RGB observations and structured unit tracking.
SMAC PySC2 VLM-Attention
@@ 3 Environment Design
We formalize our StarCraft II environment as a Partially Observable Markov Decision Process (POMDP) defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{O},P,R)$, where agents receive observations and take actions in discrete time steps.
@@ 3.4 Battle Modes and Technical Features
Figure 2: VLM-Agent Architecture integrates VLM-Attention, a knowledge-enhanced decision system, and a dynamic role assignment framework.
4 VLM-Agent Architecture
Our VLM-agent architecture integrates three key components: a self-attention mechanism for strategic unit targeting, a knowledge-enhanced decision system, and a dynamic role assignment framework. This design enables effective battlefield analysis and coordinated unit control while maintaining alignment with human decision-making processes.
4.1 VLM-Attention Mechanism
We propose VLM-Attention, an integrated system that combines skill planning and tactical decision-making through vision-language models. The mechanism processes battlefield information through two key stages to emulate human-like attention and decision-making processes.
@@ 4.2 Knowledge Integration through RAG
To enhance the VLM’s tactical understanding with domain expertise, we implement a RAG system that operates on the priority units identified by VLM-Attention. Given the priority unit set $U_{\text{priority}}\subseteq A$, we formulate the knowledge retrieval and integration process as:
@@
For example, given a priority Marine unit $u\in U_{\text{priority}}$ identified by VLM-Attention, the system retrieves structured knowledge $K(u)$ including combat statistics (DPS: 9.8 (+1.6)), counter relationships (Strong against: [Hydralisk, Immortal, Marauder]), and matchup-specific insights. The system then generates tactical guidance such as:
@@ 4.4 Complete Pipeline
We formalize our decision-making pipeline as an iterative process that maps POMDP observations to actions through a series of VLM-based transformations. At each timestep $t$, given an observation $o_{t}=(I_{t},T_{t},U_{t})$ from the environment state $s_{t}\in\mathcal{S}$, our system generates actions through the Algorithm 1.
@@
Algorithm 1 VLM-Attention Decision Pipeline for StarCraft II
@@
1:  Initialize environment and get initial observation $o_{0}=(I_{0},T_{0},U_{0})=env.reset()$
@@ 5 Experiment
Table 2: Performance evaluation of VLM-Attention across different StarCraft II scenarios, used Qwen-VL-Plus. Results show varying success rates across different unit compositions and tactical situations, with higher performance in simpler scenarios and challenges in more complex engagements.
@@
Table 4: Ablation study on vlm_attention_1 map by GPT-4-Turbo. Checkmarks indicate which components are enabled in each configuration. Results demonstrate the individual and combined effects of Role Assignment, VLM-Attention mechanism, and Retrieval-Augmented Generation (RAG) on system performance.
Role Assignment VLM-Attention RAG Win Rate (%)
@@ 5.1 Experimental Setup
We evaluate our VLM-Attention framework across 12 specialized StarCraft II micromanagement scenarios. These scenarios are carefully designed to test different aspects of tactical decision-making, ranging from basic unit control to complex multi-unit coordination with ability usage.
@@ 5.2 Component Analysis Results
Our ablation study investigates the contribution of each key component in our framework: VLM-Attention mechanism, RAG system, and Role Assignment module. The results demonstrate strong synergistic effects between these components, with the complete system achieving the highest performance (87% win rate on vlm_attention_1).
VLM-Attention proves to be the most crucial component, particularly in complex scenarios requiring sophisticated target selection and tactical positioning. When enabled alone, it improves win rates by 30-40% compared to the baseline, highlighting its effectiveness in strategic decision-making.
@@ 6 Conclusion
VLM-Attention represents a significant advancement in creating more human-like StarCraft II agents. By aligning artificial agent perception and decision-making processes with human cognition, our framework provides a more natural and effective environment for developing and testing AI systems. Future work will focus on expanding the system’s capabilities and evaluating its effectiveness in more complex scenarios.
@@ Appendix A Map Details
Our environment features a diverse collection of 21 specialized maps, systematically categorized based on player count and ability usage capabilities. These maps originate from three primary sources: SMAC-based maps redesigned from the StarCraft Multi-Agent Challenge framework, original maps specifically designed for VLM-Attention evaluation, and selected scenarios adapted from the LLM-PySC2 framework [^3] https://github.com/NKAI-Decision-Team/LLM-PySC2.
@@ A.4 Two Player Maps (With Abilities)
(c) VLM-Attention’s strategic analysis highlighting prioritized targets and optimal attack vectors
Figure 4: Visualization of VLM-Attention’s decision-making process in a complex micro-management scenario. The system processes the raw visual input (a), analyzes unit positions and states (b), and makes tactical decisions (c). In this case, VLM-Attention identified Zergling_52 (Tag: 54) as a priority target due to its strategic position at [2,1], where attacking it would maximize area-of-effect damage to nearby clustered units. This decision demonstrates the system’s ability to not only identify low-health targets (5/35 HP) but also recognize opportunities for efficient damage distribution through Colossi’s line damage mechanic. Supporting this decision, the system also identified Zergling_1 (Tag: 3) and Zergling_2 (Tag: 4) as secondary priority targets due to their threatening positions at [1,1] and [0,1] ,respectively, enabling a comprehensive control strategy that combines focus fire with positional advantage.
@@
Figure 5: Demonstration of VLM-Attention’s micro-management decision process in a Marine/Tank vs Baneling/Zergling scenario. The system implements a three-stage strategy: (1) Initial threat assessment prioritizing Banelings, (2) Coordinated unit positioning with Marines at safe distances ([0,1] and [0,2]) while maintaining focus fire capability, and (3) Sequential target elimination following the identified priority order of Banelings first, then supporting Zerglings. This showcases the system’s ability to combine Focus Fire, Formation Control, and Shield/Health Management skills in a complex tactical situation.
@@
Figure 6: Visualization of VLM-Attention’s Marine split micro-management against Siege Tanks. The system first analyzes the battlefield state (a), identifying two Siege Tanks (Tag: 10, 11, Health: 160/160) as critical threats. Then it executes a sophisticated tactical response (b) by implementing a three-way Marine split formation (3-3-2) to minimize splash damage vulnerability while maintaining effective focus fire. The coordinated execution combines multiple micro-skills: Focus Fire for optimal damage distribution, Formation Control through strategic three-way positioning, and Health Management with Medivac support positioned at [0,2]. This demonstrates the system’s ability to balance multiple tactical objectives while maintaining combat effectiveness against superior firepower.
@@
Figure 7: VLM-Attention’s multi-type unit coordination in a complex Protoss vs Terran engagement. The system first performs threat assessment (a) by identifying critical targe,ts including a low-health Viking Assault (11/125 HP), energy-rich Ghost (56 energy), and support units like Medivac. Based on this analysis, it executes a coordinated attack plan (b) where each unit is assigned optimal targets: Zealot engages the weakened Viking, Phoenix provides air superiority against Medivac, Immortal focuses on armored targets, while the Archon maintains a strategic position at [7,7] for battlefield control. The decision-making process demonstrates VLM’s understanding of unit-specific attributes (health states, energy levels, armor types) and tactical synergies in a mixed-unit combat scenario. This showcases the system’s ability to handle complex micro-management tasks involving heterogeneous unit compositions without requiring explicit training for each unit combination.

### v3

# AVA: Attentive VLM Agent for Mastering StarCraft II

@@ Abstract
We introduce Attentive VLM Agent (AVA), a multimodal StarCraft II agent that aligns artificial agent perception with the human gameplay experience. Traditional frameworks such as SMAC rely on abstract state representations that diverge significantly from human perception, limiting the ecological validity of agent behavior. Our agent addresses this limitation by incorporating RGB visual inputs and natural language observations that more closely simulate human cognitive processes during gameplay. The AVA architecture consists of three integrated components: (1) a vision-language model enhanced with specialized self-attention mechanisms for strategic unit targeting and battlefield assessment, (2) a retrieval-augmented generation system that leverages domain-specific StarCraft II knowledge to inform tactical decisions, and (3) a dynamic role-based task distribution system that enables coordinated multi-agent behavior. The experimental evaluation in our proposed AVACraft environment, which contains 21 multimodal StarCraft II scenarios, demonstrates that AVA powered by foundation models (specifically Qwen-VL and GPT-4o) can execute complex tactical maneuvers without explicit training, achieving comparable performance to traditional MARL methods that require substantial training iterations. This work establishes a foundation for developing human-aligned StarCraft II agents and advances the broader research agenda of multimodal game AI. Our implementation is available at https://github.com/camel-ai/VLM-Play-StarCraft2.
@@ 1 Introduction
We present Attentive VLM Agent (AVA), a novel StarCraft II multi-modal agent that fundamentally reimagines how AI agents perceive and interact with the game world. To evaluate the performance of AVA, we propose a multimodal StarCraft II environment, AVACraft, which makes three key innovations: First, we redesign the observation space to align with human cognitive processes, incorporating both RGB visual input and natural language descriptions. This multimodal approach enables agents to process battlefield information in a manner analogous to human players, facilitating more intuitive strategic decision-making. Second, we develop a comprehensive action space that supports complex military maneuvers, including unit targeting, formation control, and ability usage. This design allows for fine-grained tactical control while maintaining accessibility for high-level strategic planning. Third, we introduce an integrated agent architecture in AVA that combines the reasoning capabilities of VLMs with specialized mechanisms for multi-agent coordination. We incorporate self-attention mechanisms for strategic unit targeting, retrieval-augmented generation (RAG) for leveraging domain-specific knowledge, and dynamic role assignment for task distribution.
We evaluate AVA in 12 diverse micromanagement scenarios that test various aspects of tactical decision-making and unit control. Our experimental results demonstrate that VLM-based agents, utilizing models such as Qwen-VL and GPT-4o, can effectively execute complex tactical maneuvers without requiring extensive training, a significant departure from traditional MARL approaches. The primary contributions of this work include:
  • We design a multimodal StarCraft II environment (AVACraft) that aligns agent perception with human cognitive processes, featuring RGB visual inputs, natural language observations, and a comprehensive action space for complex unit control.
  • We propose an integrated VLM-based agent architecture (AVA) that combines self-attention mechanisms, retrieval-augmented generation, and dynamic role assignment for effective tactical decision-making.
  • The empirical evidence demonstrating the effectiveness of AVA in executing complex StarCraft II micromanagement tasks within AVACraft without explicit training, supported by extensive evaluation across 12 scenarios.
Our work establishes a foundation for developing more human-aligned StarCraft II agents and contributes to the broader goal of creating AI systems that can understand and interact with complex game environments in ways that more closely mirror human cognition.
@@ 2 Related Work
3 AVACraft Environment Design
Figure 1: Observation Space of AVACraft environment.
Table 1: Comparison of StarCraft II environments. While SMAC simplifies the interface for MARL training and PySC2 provides comprehensive API access, AVACraft adapts PySC2 for multimodal decision-making through RGB observations and structured unit tracking.
SMAC PySC2 AVACraft
@@ 3 AVACraft Environment Design
We formalize our AVACraft environment as a Partially Observable Markov Decision Process (POMDP) defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{O},P,R)$, where agents receive observations and take actions in discrete time steps.
@@ 3.4 Battle Modes and Technical Features
Figure 2: VLM-Agent Architecture integrates AVA, a knowledge-enhanced decision system, and a dynamic role assignment framework.
4 Attentive VLM Agent Architecture
Our Attentive VLM Agent (AVA) architecture integrates three key components: a self-attention mechanism for strategic unit targeting, a knowledge-enhanced decision system, and a dynamic role assignment framework. This design enables effective battlefield analysis and coordinated unit control while maintaining alignment with human decision-making processes.
4.1 AVA Mechanism
We propose AVA, an integrated system that combines skill planning and tactical decision-making through vision-language models. The mechanism processes battlefield information through two key stages to emulate human-like attention and decision-making processes.
@@ 4.2 Knowledge Integration through RAG
To enhance the VLM’s tactical understanding with domain expertise, we implement a RAG system that operates on the priority units identified by AVA. Given the priority unit set $U_{\text{priority}}\subseteq A$, we formulate the knowledge retrieval and integration process as:
@@
For example, given a priority Marine unit $u\in U_{\text{priority}}$ identified by AVA, the system retrieves structured knowledge $K(u)$ including combat statistics (DPS: 9.8 (+1.6)), counter relationships (Strong against: [Hydralisk, Immortal, Marauder]), and matchup-specific insights. The system then generates tactical guidance such as:
@@ 4.4 Complete Pipeline
We formalize our decision-making pipeline as an iterative process that maps POMDP observations to actions through a series of VLM-based transformations. At each timestep $t$, given an observation $o_{t}=(I_{t},T_{t},U_{t})$ from the AVACraft environment state $s_{t}\in\mathcal{S}$, our system generates actions through the Algorithm 1.
@@
Algorithm 1 AVA Decision Pipeline for AVACraft
@@
1:  Initialize AVACraft environment and get initial observation $o_{0}=(I_{0},T_{0},U_{0})=env.reset()$
@@ 5 Experiment
Table 2: Performance evaluation of AVA across different StarCraft II scenarios, used Qwen-VL-Plus. Results show varying success rates across different unit compositions and tactical situations, with higher performance in simpler scenarios and challenges in more complex engagements.
@@
Table 4: Ablation study on vlm_attention_1 map by GPT-4-Turbo. Checkmarks indicate which components are enabled in each configuration. Results demonstrate the individual and combined effects of Role Assignment, AVA mechanism, and Retrieval-Augmented Generation (RAG) on system performance.
Role Assignment AVA RAG Win Rate (%)
@@ 5.1 Experimental Setup
We evaluate our AVA framework across 12 specialized StarCraft II micromanagement scenarios. These scenarios are carefully designed to test different aspects of tactical decision-making, ranging from basic unit control to complex multi-unit coordination with ability usage.
@@ 5.2 Component Analysis Results
Our ablation study investigates the contribution of each key component in our framework: AVA mechanism, RAG system, and Role Assignment module. The results demonstrate strong synergistic effects between these components, with the complete system achieving the highest performance (87% win rate on vlm_attention_1).
AVA proves to be the most crucial component, particularly in complex scenarios requiring sophisticated target selection and tactical positioning. When enabled alone, it improves win rates by 30-40% compared to the baseline, highlighting its effectiveness in strategic decision-making.
@@ 6 Conclusion
AVA represents a significant advancement in creating more human-like StarCraft II agents. By aligning artificial agent perception and decision-making processes with human cognition, our framework provides a more natural and effective environment for developing and testing AI systems. Future work will focus on expanding the system’s capabilities and evaluating its effectiveness in more complex scenarios.
@@ Appendix A Map Details
Our AVACraft environment features a diverse collection of 21 specialized maps, systematically categorized based on player count and ability usage capabilities. These maps originate from three primary sources: SMAC-based maps redesigned from the StarCraft Multi-Agent Challenge framework, original maps specifically designed for AVA evaluation, and selected scenarios adapted from the LLM-PySC2 framework [^3] https://github.com/NKAI-Decision-Team/LLM-PySC2.
@@ A.4 Two Player Maps (With Abilities)
(c) AVA’s strategic analysis highlighting prioritized targets and optimal attack vectors
Figure 4: Visualization of AVA’s decision-making process in a complex micro-management scenario. The system processes the raw visual input (a), analyzes unit positions and states (b), and makes tactical decisions (c). In this case, AVA identified Zergling_52 (Tag: 54) as a priority target due to its strategic position at [2,1], where attacking it would maximize area-of-effect damage to nearby clustered units. This decision demonstrates the system’s ability to not only identify low-health targets (5/35 HP) but also recognize opportunities for efficient damage distribution through Colossi’s line damage mechanic. Supporting this decision, the system also identified Zergling_1 (Tag: 3) and Zergling_2 (Tag: 4) as secondary priority targets due to their threatening positions at [1,1] and [0,1] ,respectively, enabling a comprehensive control strategy that combines focus fire with positional advantage.
@@
Figure 5: Demonstration of AVA’s micro-management decision process in a Marine/Tank vs Baneling/Zergling scenario. The system implements a three-stage strategy: (1) Initial threat assessment prioritizing Banelings, (2) Coordinated unit positioning with Marines at safe distances ([0,1] and [0,2]) while maintaining focus fire capability, and (3) Sequential target elimination following the identified priority order of Banelings first, then supporting Zerglings. This showcases the system’s ability to combine Focus Fire, Formation Control, and Shield/Health Management skills in a complex tactical situation.
@@
Figure 6: Visualization of AVA’s Marine split micro-management against Siege Tanks. The system first analyzes the battlefield state (a), identifying two Siege Tanks (Tag: 10, 11, Health: 160/160) as critical threats. Then it executes a sophisticated tactical response (b) by implementing a three-way Marine split formation (3-3-2) to minimize splash damage vulnerability while maintaining effective focus fire. The coordinated execution combines multiple micro-skills: Focus Fire for optimal damage distribution, Formation Control through strategic three-way positioning, and Health Management with Medivac support positioned at [0,2]. This demonstrates the system’s ability to balance multiple tactical objectives while maintaining combat effectiveness against superior firepower.
@@
Figure 7: AVA’s multi-type unit coordination in a complex Protoss vs Terran engagement. The system first performs threat assessment (a) by identifying critical targe,ts including a low-health Viking Assault (11/125 HP), energy-rich Ghost (56 energy), and support units like Medivac. Based on this analysis, it executes a coordinated attack plan (b) where each unit is assigned optimal targets: Zealot engages the weakened Viking, Phoenix provides air superiority against Medivac, Immortal focuses on armored targets, while the Archon maintains a strategic position at [7,7] for battlefield control. The decision-making process demonstrates VLM’s understanding of unit-specific attributes (health states, energy levels, armor types) and tactical synergies in a mixed-unit combat scenario. This showcases the system’s ability to handle complex micro-management tasks involving heterogeneous unit compositions without requiring explicit training for each unit combination.